Complete the table using the given values of x and the equation -2x – y = 1. Show your work.
X Y
-1
0
1
2
3
4

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Complete the table using the given values of x and the equation -2x – y = 1. Show your work.
X Y
-1
0
1
2
3
4
===============
y = -2x - 1
Sub the values for x, find y.

Answer by stanbon(75887)

Complete the table using the given values of x and the equation -2x – y = 1. Show your work.
X..... Y = -2x-1
-1 ... Y = -2*-1-1 = 1
0 .... Y = -2*0-1 = -1
1 .... y = -2*1-1 = -3
etc.
Cheers,
Stan H.
=============
2
3
4